# Learning Domain-Invariant Features for Out-of-Context News Detection

### v1

@@ 4.2. Domain-Invariant Feature Learning
(8) $\mathcal{L}_{CE}=-\mathbf{E}_{y\sim Y}\left[y\log(\hat{y})+(1-y)\log(1-\hat{y})\right].$
@@
(9) $\mathcal{L}=\frac{1}{2}\lambda_{CE}(\mathcal{L}_{CE}^{S}+\mathcal{L}_{CE+}^{S})+\frac{1}{2}\lambda_{ctr}(\mathcal{L}_{ctr}^{S}+\mathcal{L}_{ctr}^{T})+\lambda_{MMD}\mathrm{MMD},$
@@
4.3. Test-Time Adaptation
Domain-invariant features help the model to generalize to the target domain. In this section, we introduce test-time adaptation which leverages target domain statistics to further adapt the model to the target domain during evaluation.
Prior work (Li et al., 2016) observes that the statistics of the Batch Normalization (BN) layer contain domain-specific information. Concretely, they find that running the test set through the trained model achieves deep adaptation effects on the target domain. The reason is that this process allows BN statistics to keep track of the target domain statistics, which is conducive to the domain adaptation. As such, we introduce BN layers into the classifier.
@@ 4.3. Test-Time Adaptation
(10) $\hat{\mu}_{k+1}=(1-\rho)\hat{\mu}_{k}+\rho\mu_{k},\quad\hat{\sigma}^{2}_{k+1}=(1-\rho)\hat{\sigma}^{2}_{k}+\rho\sigma^{2}_{k}$
(11) $\mu_{k}=\frac{1}{b}\sum_{i}x_{i},\quad\sigma^{2}_{k}=\frac{1}{b}\sum_{i}(x_{i}-\mu_{k})^{2}$
Here, $k$ denotes the $k$-th training step, and $\rho$ is the momentum with a default value of 0.1. BN statistics are then fixed (denoted as $\mu$ and $\sigma^{2}$) and used for normalization when the model is set to evaluation mode (pyt, 2024). During evaluation, we apply test-time adaptation to calibrate the BN statistics, i.e. the mean ($\mu$) and variance ($\sigma^{2}$). The BN layer is formulated as follows:
(12) $x_{out}=\frac{x_{in}-\mu}{\sqrt{\sigma^{2}+\epsilon}}*\gamma+\beta$
where $x_{in}$ is the BN layer input, $x_{out}$ is the output. $\gamma$ and $\beta$ are BN learnable parameters. $\epsilon$ is added for numerical stability.
During model training, since only the source domain data is passed through the classifier, BN statistics are computed based on the source domain statistics. During model evaluation, the target domain data is passed through the classifier. To integrate target domain statistics into the BN statistics, after each training epoch, we pass the unlabeled test set through the classifier. After that, we set the classifier to evaluation mode and evaluate it on the test set. In this way, the running estimates would keep the target domain statistics as well into record, which is helpful for the domain adaptation.
5. Experimental Design
5.1. Datasets
@@
Twitter-COMMs (Biamby et al., 2022): This dataset is collected from Twitter solely on three topics: Covid-19 (Cv), Climate Change (Cl) and Military Vehicles (M). Each original tweet is represented by an image and its associated caption (text). Given a caption, the authors create out-of-context news by either randomly choosing an image (random negative) or by choosing the image whose original associated text is semantically the closest (hard negative). After running the script provided by the authors to download tweets and removing tweets that has an unavailable image, we obtain 2,143,934 items (Cv: 1,387,043, Cl: 512,490, M: 244,401) in the training set, and 22,082 items (Cv: 6,456, Cl: 8,488, M: 7,138) in the test set. The labels have a balanced distribution. Hence, we address domain adaptation among different news topics in Twitter-COMMs.
NewsCLIPpings (Luo et al., 2021): This dataset is derived from the VisualNews dataset (Liu et al., 2020), which is a benchmark news image captioning dataset. The news is collected from four agencies, namely Guardian (G), BBC (B), Washington Post (W) and USA Today (U). The NewsCLIPpings dataset is originally partitioned into four splits in which ways the out-of-context news is constructed. However, the analysis on the underlying VisualNews dataset (Liu et al., 2020) reveals that there are inter correlations among data from different news agencies (e.g. , the image captioning model trained on the USA Today data achieves 3.7 in CIDEr score (Vedantam et al., 2015) when tested on the USA Today test set, yet only achieves 0.6 on the Guardian test set. ). Since NewsCLIPping is derived from VisualNews, we re-partition the dataset according to the news agencies. After downloading the dataset, we obtain a total number of 1,182,900 items (B: 205,968, G: 567,012, U: 228,393, W: 181,527) in the training set, and 124,168 items (B: 22,074, G: 59,163, U: 24,243, W: 18,688) in the test set. The labels have a balanced distribution. Hence, we address domain adaptation among different news agencies in NewsCLIPpings.
5.2. Implementation Details
Table 1. Evaluation of domain adaptation performance on TwitterCOMMs and NewsCLIPpings. Cv, Cl and M stand for covid-19, climate change and military vehicles respectively. B, G, U and W stand for BBC, Guardian, USA Today and Washington Post respectively. In X $\rightarrow$ Y, X denotes the source domain, Y denotes the target domain. The best performance is in bold text. The second best performance is underlined.
@@ 5.2. Implementation Details
Multimodal Feature Encoder Selection: To find the optimal MLLM feature extractor, we experiment with (i) BLIP-2 (Li et al., 2023b) multimodal and (ii) unimodal feature extractor, (iii) CLIP (Radford et al., 2021) (only has unimodal feature extractors) and (iv) ALBEF (Li et al., 2021b) multimodal feature extractor. For (ii) and (iii), we experiment with vector concatenation and element-wise multiplication to obtain the multimodal feature vector.
We then train a one-layer linear classifier on the four features respectively and evaluate the performances. Experimental results show that BLIP-2 multimodal features gives the best performance (in terms of accuracy) on both the training and test sets. Note, we also evaluate the zero-shot performance of the BLIP-2 Image Text Matching head which only outputs a similarity score rather than the feature vector. We find it underperforms (around 0.5 in accuracy on both training and test set) using BLIP-2 multimodal feature directly.
Architecture: We use BLIP-2’s blip2_feature_extractor with the multimodal mode to embed the image and text, and obtain the multimodal embedding x of size 768. The projection head is composed of two linear layers with 768 and 500 (experimented with {300, 500, 768}) neurons respectively. The classifier is composed of three linear layers with 768, 768 and 2 neurons respectively, and two batch normalization layers following the first two linear layers. We use Tanh() as the activation function. We apply dropout both before the first linear layer and the third linear layer. The dropout rate is set to 0.2 (we experimented with {0.2, 0.5} and without dropout).
Training: We perform training using the Adam optimizer with a batch size of 256 for 20 epochs. The early stopping epoch is set to 5. The learning rate is set to 2e-4 for the Twitter-COMMs dataset and 1e-4 for the NewsCLIPpings dataset. For the loss function, $\lambda_{MMD}$ is set to 1 and both $\lambda_{CE}$ and $\lambda_{ctr}$ are set to 0.5.
[…]
5.3. Baseline Models
Considering that no prior work has looked into domain adaptive out-of-context news detection, we select SOTA baseline models from a similar task: domain adaptive fake news detection. For fair comparison, we use BLIP-2’s blip2_feature_extractor with the multimodal mode to encode the image and text into a multimodal feature vector.
@@ 5.3. Baseline Models
EANN (Wang et al., 2018) is a multimodal fake news detection model that uses adversarial learning to derive event-invariant features which benefit fake news detection on newly arrived events. Since the original EANN implementation encodes the image and text modality separately, we modify it to accommodate multimodal inputs.
@@
MDA-WS (Li et al., 2021a) is an early fake news detection framework based on multi-source domain adaptation and weak supervision. It uses adversarial learning to learn the domain-invariant feature, and incorporates prior knowledge to assign pseudo labels to target domain news for weak supervision.
@@
CANMD (Yue et al., 2022) is a domain adaptive fake news detection model. It firstly pre-trains the model on the source training data and then uses the pseudo labels generated from it on the target training data to adapt the model. During adaptation, it aims to reduce intra-class discrepancy and enlarge inter-class discrepancy.
@@
REAL-FND (Mosallanezhad et al., 2022) is a reinforcement-learning (RL) based domain adaptive fake news detection model. It first obtains a reasonable news representation and then trains the fake news and domain classifiers. Using the representation as the state and the classifiers’ losses as the reward, it then trains an RL agent that learns a domain-invariant news representation.
6. Result and Analysis
@@
Based on the above experimental setup, we evaluate our proposed ConDA-TTA and investigate the contribution of each model component afterwards. We also include TSNE visualization and parameter sensitivity analysis.
6.1. Experimental Result
Table 2. Evaluation results on ablating $\mathcal{L}_{ctr}$, MMD and TTA of ConDA-TTA. Cv, Cl and M stand for covid-19, climate change and military vehicles respectively. B, G, U and W stand for BBC, Guardian, USA Today and Washington Post respectively. In X $\rightarrow$ Y, X denotes the source domain, Y denotes the target domain. The best performance is in bold text. The second best performance is underlined.
@@ 6.1. Experimental Result
Table 1 shows the evaluation performance of ConDA-TTA and the baseline models. We use F1 and accuracy (Acc) as the evaluation metrics.
Naive Baseline and Upper Bound: We first present our observations on the naive baseline (source only in Table 1) and the supervised training (source and target in Table 1) results. In our experiments, we find that using BLIP-2’s multimodal representation with a plain linear classifier already achieves strong performances (around 70 - 80% zero-shot accuracy under most domain adaptation scenarios) on this task. Nevertheless, we observe around 2-10% performance gap between the naive baseline and the supervised training result. Considering that BLIP-2 has 2.7b parameters, it is not trivial to apply fine-tuning to improve its domain adaptation performance. This suggests the need for developing domain adaptation techniques on top of it to mitigate the gap.
For the news topics in Twitter-COMMs, adapting to Military Vehicles (see Section 5.1) has the largest domain adaptation gap (10.6% in F1 and 8.32% in accuracy), whilst adapting to Climate Change and Covid-19 has a relatively smaller domain adaptation gap (2.99% in F1 and 2.65% in accuracy, 4.58% in F1 and 3.71% in accuracy). We conjecture that the differences are potentially caused by the different news styles. We manually check a batch of 10 randomly sampled news from each topic. We find that Climate Change and Covid-19 tend to include images with more abundant information, including things such as embedded texts, numbers, and even news screenshots. However, for the Military Vehicles domain, the news’ image usually contains one clear subject: the military vehicle itself. At the same time, it requires more specialized knowledge to determine whether the image pairs the text within this topic.
For the news agencies in NewsCLIPpings, adapting to BBC and USA Today has the largest domain adaptation gap (6.61% in F1 and 6.81% in accuracy, 2.93% in F1 and 4.43% in accuracy), while adapting to the Guardian and the Washington Post has a relatively smaller gap. Although the gap is not consistently significant, it can affect heavily if deployed in a news platforms where a large amount of news is posted everyday. Thus, we believe that leveraging these powerful MLLM feature encoders and further mitigating its domain adaptation gap with a small training cost are crucial to achieve promising performance on unlabeled, newly emergent domains.
Baseline Comparison: We next compare our model’s performance with the baseline models. On both datasets, our proposed ConDA-TTA outperforms the baseline models across most metrics. Notably, on the Twitter-COMMs dataset, ConDA-TTA outperforms the best baseline model by 2.93% in F1 and by 2.08% in accuracy when the source domains are Covid-19 (Cv) and Climate Change (Cl), and the target domain is Military Vehicles (M). Specifically, we observe negative transfer for CANMD under this setting. Moreover, ConDA-TTA achieves 80.80% on both F1 and accuracy when the source domains are Covid-19 and Military Vehicles, and the target domain is Climate Change, surpassing the baseline performances by 0.93% and 0.27% respectively. Additionally, it achieves 80.06% and 80.08% on F1 and accuracy when the source domains are Climate Change and Military Vehicles, and the target domain is Covid-19.
On the NewsCLIPpings dataset, when the source domains are USA Today (U) and Washington Post (W) and the target domain is BBC (B), our model outperforms the best baseline model by 1.84% in accuracy and 1.82% F1 accuracy. When the source domains are BBC (B) and Guardian (G) and the target domain is USA Today (U), our model outperforms the best baseline model by 0.82% in F1 and 0.84% in accuracy. When the source domains are USA Today and Washington Post, and the target domain is Guardian, our model achieves the second best performance in both metrics. However, we also observe that when the source domains are BBC and Guardian, and target domain is Washington Post, all models see slightly negative transfer in terms of accuracy. This is in-line with the observations (Section  5.1) on the underlying dataset (Liu et al., 2020), where the authors find that training their image captioning model on either the BBC or Guardian results in a comparatively low CIDEr score on the Washington Post. As such, we conjecture that the information contained in the BBC and Guardian is not enough for enhancing the model’s domain adaptation performance on the Washington Post.
@@ 6.2. Ablation Study
We next present an ablation study on the different components in ConDA-TTA. Specifically, we ablate $\mathcal{L}_{ctr}$, MMD and TTA from ConDA-TTA. The evaluation results are summarized in Table 2.
@@
On the Twitter-COMMs dataset, we observe that ablating MMD causes the biggest performance drop in all three domain adaptation settings, respectively: 3.41% (F1) and 3.42% (Acc), 0.33% (F1) and 0.32% (Acc), 0.92% (F1) and 0.91% (Acc). This suggests that MMD is the most important component in ConDA-TTA. Additionally, ablating the contrastive loss results in the second biggest performance decrease when the target domain is Military Vehicles and Climate Change, respectively: 1.19% (F1) and 1.22% (Acc), 0.21% (F1) and 0.21% (Acc). However, we see a marginal performance increase (0.04% in F1 and 0.03% in Acc) when the target domain is Covid-19. Lastly, ablating TTA causes a 0.68% (F1) and 0.75% (Acc) performance drop when the target domain is Military Vehicles, and causes a 0.07% performance drop in both F1 and Acc when the target domain is Climate Change. When the target domain is Covid-19, we again see a slight increase (0.46% in F1 and 0.50% in Acc) in performance. Although ConDA-TTA (as a whole) improves the domain adaptation performance on Covid-19 (Section 6.1), not every component of it contributes under this setting.
@@
On the NewsCLIPpings dataset, we observe similar degradation when ablating these three components when the target domain is the BBC, Guardian and USA Today. Specifically, when ablating TTA, the performance decreases (1.22% in F1 ad 1.53% in Acc) most significantly on BBC, while less significantly on Guardian (0.02% in F1 and 0.18% in Acc) and USA Today (0.02% in F1 and 0.07% in Acc). When ablating MMD, F1 drops by 0.04%, 0.06% and 0.84%, and the Acc drops by 0.04%, 0.06% and 0.86%. When ablating the contrastive loss, the F1 drops by 0.62%, 0.63% and 0.5%, and the Acc drops by 0.53%, 0.58% and 0.51%. However, when adapting to Washington Post, ablating the contrastive loss, MMD and TTA all bring slight performance increase (0.04% - 0.72%). We conjecture that this observation echos previous analysis (Section 6.1) that the information contained in BBC and Guardian is not enough for assisting the domain adaptation on Washington Post, making each component less effective.
In summary, the above results confirm the efficacy of the contrastive loss, MMD and TTA in most domain adaptation scenarios on both datasets. Concretely, MMD tends to contribute the most to the domain adaptation on Twitter-COMMs; TTA tends to contribute the most on NewsCLIPpings. However, we note that, in some cases, the techniques are less effective on the NewsCLIPpings dataset. We believe this is due to that under some settings, the source domain contains less abundant information for adaptation to the target domain (Section 6.1).
6.3. Visualization

### v2

@@ 4.2 Contrastive Domain Adaptation
$\mathcal{L}_{CE}=-\mathbf{E}_{y\sim Y}\left[y\log(\hat{y})+(1-y)\log(1-\hat{y})\right].$ (8)
@@
$\mathcal{L}=$ $\frac{1}{2}\lambda_{CE}(\mathcal{L}_{CE}^{S}+\mathcal{L}_{CE+}^{S})+\frac{1}{2}\lambda_{ctr}(\mathcal{L}_{ctr}^{S}+\mathcal{L}_{ctr}^{T})$ (9)
$+\lambda_{MMD}\mathrm{MMD},$
@@
4.3 Test-Time Adaptation
Domain-invariant features help the model generalize to the target domain. In this section, we introduce test-time adaptation which leverages target domain statistics to further adapt the model to the target domain during evaluation.
Prior work Li et al. (2016) observes that the statistics of the Batch Normalization (BN) layer contain domain-specific information. Concretely, they find that running test set through the trained model achieves deep adaptation effects on the target domain. The reason is that this process allows BN statistics to keep track of the target domain statistics, which is conducive to the domain adaptation.
@@ 4.3 Test-Time Adaptation
$\hat{\mu}_{k+1}=(1-\rho)\hat{\mu}_{k}+\rho\mu_{k},\quad\hat{\sigma}^{2}_{k+1}=(1-\rho)\hat{\sigma}^{2}_{k}+\rho\sigma^{2}_{k},$ (10)
$\mu_{k}=\frac{1}{b}\sum_{i}x_{i},\quad\sigma^{2}_{k}=\frac{1}{b}\sum_{i}(x_{i}-\mu_{k})^{2}.$ (11)
Here, $k$ denotes the $k$-th training step, and $\rho$ is the momentum with a default value of 0.1. $x_{i}$ denotes the input to BN. BN statistics are then fixed (denoted as $\mu$ and $\sigma^{2}$) and used for normalization when the model is set to evaluation mode pyt (2024).
As such, we add BN layers into the classifier. To integrate target domain statistics into BN’s memory, we first pass the unlabeled target domain test data through the classifier before setting it to the evaluation mode. After that, we set the classifier to evaluation mode and evaluate it on the test set. In this way, the running estimates would keep the target domain statistics as well into record, which is helpful for the domain adaptation.
5 Experimental Design
5.1 Datasets
@@
Twitter-COMMs Biamby et al. (2022): This dataset is collected from Twitter solely on three topics: Covid-19 (Cv), Climate Change (Cl) and Military Vehicles (M). After downloading the tweets and removing the tweets that has an unavailable image, we obtain 2,143,934 items (Cv: 1,387,043, Cl: 512,490, M: 244,401) in the training set, and 22,082 items (Cv: 6,456, Cl: 8,488, M: 7,138) in the test set. The labels are balanced.
NewsCLIPpings Luo et al. (2021): This dataset is derived from the VisualNews dataset (Liu et al., 2020), a benchmark news image captioning dataset. The news is collected from four agencies, namely Guardian (G), BBC (B), Washington Post (W) and USA Today (U). After downloading the dataset, we obtain a total number of 1,182,900 items (B: 205,968, G: 567,012, U: 228,393, W: 181,527) in the training set, and 124,168 items (B: 22,074, G: 59,163, U: 24,243, W: 18,688) in the test set. The labels are balanced.
5.2 Implementation Details
Augmentation: In our approach, we apply Gaussian blur to the image while keeping the text unaltered to generate the augmentation of the anchor item. More details and experimental results can be found in Appendix A.1.
Architecture: We use BLIP-2’s multimodal feature extractor to embed the image and text, and obtain the embedding x of size 768. Details of multimodal feature encoder selection can be found in Appendix A.1. The projection head is composed of two linear layers with 768 and 500 (experimented with {300, 500, 768}) neurons respectively. The classifier is composed of three linear layers with 768, 768 and 2 neurons respectively, and two batch normalization layers following the first two linear layers. We use Tanh() as the activation function. We apply dropout both before the first and the third linear layer. The dropout rate is set to 0.2 (we experiment with {0.2, 0.5} and without dropout).
Training: We perform training using the Adam optimizer with a batch size of 256 for 20 epochs. The early stopping epoch is set to 5. The learning rate is set to 2e-4 for Twitter-COMMs and 1e-4 for NewsCLIPpings. For the loss function, $\lambda_{MMD}$ is set to 1 and both $\lambda_{CE}$ and $\lambda_{ctr}$ are set to 0.5.
Source and Target Domain Partition: We treat news toics as individual domains in Twitter-COMMs. We use one topic as the target domain and the remaining two topics as the source domain. We treat news agencies as individual domains in NewsCLIPpings. Based on an observation (Liu et al., 2020) on the correlation of these four news agencies and our pilot experiments, we use two news agencies of the same country as the target domain and the rest two as the source domain respectively.
5.3 Baseline Models
Table 1: Evaluation of domain adaptation performance on TwitterCOMMs and NewsCLIPpings. Cv, Cl and M stand for covid-19, climate change and military vehicles respectively. B, G, U and W stand for BBC, Guardian, USA Today and Washington Post respectively. In X $\rightarrow$ Y, X denotes the source domain, Y denotes the target domain. The best performance is in bold text. The second best performance is underlined.
@@ 5.3 Baseline Models
Considering that no prior work has looked into domain adaptive out-of-context news detection, we select SOTA baseline models from a similar task: domain adaptive fake news detection. For fair comparison, we use the same multimodal feature encoder for all baselines.
@@
EANN Wang et al. (2018) is a multimodal fake news detection model that uses adversarial learning to derive event-invariant features which benefit fake news detection on newly arrived events.
@@
MDA-WS Li et al. (2021b) uses adversarial learning to learn the domain-invariant feature, and incorporates prior knowledge to assign pseudo labels to target domain news for weak supervision.
@@
CANMD Yue et al. (2022) firstly pre-trains the model on the source training data and then generates pseudo labels for the target training data to adapt the model.
@@
REAL-FND Mosallanezhad et al. (2022) uses news representation as the state and the classifiers’ losses as the reward, and trains an RL agent that learns a domain-invariant news representation.
CADA Li et al. (2023a) proposes class-based adversarial domain adaptive framework to achieve fine-grained alignment.
6 Result and Analysis
@@
Based on the above experimental setup, we evaluate our proposed ConDA-TTA and investigate the contribution of each model component afterwards. We also include TSNE visualization and parameter sensitivity analysis (Appendix A.3).
6.1 Experimental Result
Table 1 shows the evaluation performance of ConDA-TTA and the baseline models. We use F1 (%) and accuracy (Acc%) as the evaluation metrics as in Luo et al. (2021) and Biamby et al. (2022).
Naive Baseline and Upper Bound: We first present our observations on the naive baseline (source only) and the supervised training (source and target) results. In our experiments, we find that using BLIP-2’s multimodal representation with a plain linear classifier gives competitive performances. However, we observe around 2-10% performance gap between the naive baseline and the supervised training. This suggests the need for developing domain adaptation techniques to mitigate the gap.
The domain adaptation gaps of different settings vary. For news topics in Twitter-COMMs, adapting to Military Vehicles (M) has the largest domain adaptation gap (10.6% in F1), whilst adapting to Climate Change (Cl) and Covid-19 (Cv) have relatively smaller domain adaptation gaps. We conjecture that the differences are potentially caused by the different news styles. We manually check a batch of 10 randomly sampled news from each topic. We find that Cl and Cv tend to include images that have more abundant information, containing such as embedded texts, numbers, and even news screenshots. However, in M, the news image usually contains one clear subject: the vehicle itself. At the same time, it requires more specialized knowledge to determine whether the image pairs the text within this topic.
For news agencies in NewsCLIPpings, adapting to BBC (B) and USA Today (U) have a larger domain adaptation gap (6.81% and 4.43% in accuracy), while adapting to the Guardian (G) and the Washington Post (W) have a relatively smaller gap. Although the gap is not consistently large, it could cause a big difference on news platforms where a large amount of news is posted everyday. Thus, we believe that it is crucial to leverage MLLM feature encoders and further mitigating its domain adaptation gap with a small training cost.
Baseline Comparison: We next compare our model’s performance with the baseline models. On both datasets, our proposed ConDA-TTA outperforms the baseline models across most metrics. Notably, on Twitter-COMMs, ConDA-TTA outperforms the best baseline model by 2.93% in F1 when the source domains are Covid-19 (Cv) and Climate Change (Cl), and the target domain is Military Vehicles (M).
On NewsCLIPpings, when the target domain is BBC (B) and USA Today (U), our model outperforms the baselines. When the target domain is Guardian (G), our model achieves the second best performance. However, we observe that when the target domain is the Washington Post (W), all models see slightly negative transfer. This is in line with the observations on the underlying dataset Liu et al. (2020), where the authors find that training their image captioning model on either the B or G results in a comparatively low CIDEr score (Vedantam et al., 2015) on the W. As such, we conjecture that the information contained in B and G is not enough for enhancing the model’s domain adaptation performance on W.
Table 2: Evaluation results on ablating $\mathcal{L}_{ctr}$, MMD and TTA of ConDA-TTA.
@@ 6.2 Ablation Study
We next present an ablation study on different components in ConDA-TTA. Specifically, we ablate $\mathcal{L}_{ctr}$, MMD and TTA from ConDA-TTA. The evaluation results are summarized in Table 2.
@@
On Twitter-COMMs, we observe that ablating MMD causes the biggest performance drop in all three domain adaptation settings. Additionally, ablating $\mathcal{L}_{ctr}$ and TTA both result in performance decrease when the target domain is Military Vehicles (M) and Climate Change (Cl). However, we see a marginal performance increase when the target domain is Covid-19 (Cv).
@@
On NewsCLIPpings, we observe similar degradation when ablating these three components for target domain being BBC (B), Guardian (G) and USA Today (U). Specifically, when ablating TTA, the performance decreases most significantly on B, and less significantly on G and U. However, when adapting to Washington Post (W), ablating $\mathcal{L}_{ctr}$, MMD and TTA all bring slight performance increase. We assume that this observation echos the previous analysis (Section 6.1) that the information contained in BBC and Guardian is not enough for helping the domain adaptation to Washington Post, therefore making each component less effective.
In summary, the above results confirm the efficacy of $\mathcal{L}_{ctr}$, MMD and TTA under most domain adaptation scenarios on both datasets. Concretely, MMD tends to contribute the most on Twitter-COMMs; TTA tends to contribute the most on NewsCLIPpings. It is worth noting that news in Twitter-COMMs are more colloquial and informal, leading writing styles vary much across different topics. We conclude that MMD is better at capturing domain-invariant features in this scenario.
6.3 Visualization